OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities: 2x - 3y greater than or equal to 12 x + y less than or equal to 7 (-9, -1) (3, -6) (12, -3) (6, 9) Two boats leave from the same dock traveling in opposite directions. The first boat is traveling 25 mph and the second is traveling 29 mph. How long will it take before they are 162 miles apart? 1 hour 2 hours 3 hours 4 hours

OpenStudy (anonymous):

Are those points the one that came with the question?

OpenStudy (anonymous):

do u know how to solve simultaneous equations?

OpenStudy (anonymous):

2x - 3y >= 12 x + y <= 7

OpenStudy (anonymous):

multiply both sides of the 2nd equation by -2 and add the 2 equations , know how to do this?

OpenStudy (anonymous):

im really confused on how to even do these questions. kx2bay

OpenStudy (anonymous):

just try , I gave you the first step above, do you know how to multiply ?

OpenStudy (anonymous):

-2(x + y) >= -2*7 .... notice when multitplying by -2 the sign <= changed to >= -2x - 2y >= -14

OpenStudy (anonymous):

now add the 2 equations 2x - 3y >= 12 -2x - 2y >= -14 ========== you get -5y >= -2 now divide by -5 both sides (the inequality sign will change) you'll get y <= 2/5 ******************** now can you find x ?

OpenStudy (anonymous):

Oh wow, makes sense now! we eliminated the x and we got y so now we plug the y into one of the equations to get x right?

OpenStudy (anonymous):

yep, can you do that?

OpenStudy (anonymous):

let y = 2/5 and substitute it in either of the 2 equations, perhaps use the shorter one x + y <= 7 what will x be then?

OpenStudy (anonymous):

x = -2/5 + 7 is what i got, i cant think of anything else to do to it.

OpenStudy (anonymous):

x <= 33/5

OpenStudy (anonymous):

so now you have both x and y x <= 33/5 , and y <= 2/5 and if you said x = 33/5 and y = 2/5 and substituted these values in any of the 2 equations you'll get the equation 1 is equal to 12 and equation 2 is equal to 7

OpenStudy (anonymous):

the answer is (-9, -1)

OpenStudy (anonymous):

with the second question about the boats... are you there?

OpenStudy (anonymous):

do you understand what it says?

OpenStudy (anonymous):

the formula for speed is speed = distance / time

OpenStudy (anonymous):

one boat travels left and passes x miles after some time let's call it "t" the other boat travels right and passes a different distance y for the same time "t"

OpenStudy (anonymous):

it is given that the distance between the 2 boats after the time "t" is 162 miles so, your first equation is x + y = 162 now using the law of speed can you make the second formula ?

OpenStudy (anonymous):

25 = x / t, or t = x / 25 and for the other boat 29 = y / t or t = y / 29 so... x / 25 = y / 29 cross multiply , find x in terms of y and substitute in the first formula above, then solve for the other boat

OpenStudy (anonymous):

got it????

OpenStudy (anonymous):

x = 75 miles and y = 87 miles now using one of the TIME formulas above, find "t" in hours? use t = x/25 , you know x from above, what is t ?

OpenStudy (anonymous):

ok... good luck sandy :)